# Text-guided HuBERT: Self-Supervised Speech Pre-training via Generative Adversarial Networks

### v4

@@ II-B GAN-based Tokenizer
To leverage additional text data and improve the phonetic information of the pre-training, we utilize the GANs [17] as a bridge between unpaired speech and text, to discretize the speech waveform into phoneme-like pseudo-label sequences in an unsupervised manner. As shown in the right part of Figure 1, the generator consumes speech, while the input of the discriminator is the real phoneme sequences of unpaired text. During adversarial training, the discriminator is trained to distinguish the output of the generator and real phoneme sequences. The goal of the generator is to output phoneme sequences that cannot be distinguished by the discriminator.
@@ II-C Multi-Unit and Multi-Layer Supervision Objective
Our model architecture design mainly follows the HuBERT-base model, consisting of a convolutional feature encoder and a transformer encoder. The convolutional encoder is composed of seven temporal convolution blocks, which have 512 channels with strides $(5,2,2,2,2,2,2)$ and kernel widths $(10,3,3,3,3,2,2)$, resulting in a 20ms frame rate for audio sampled at 16kHz. The transformer encoder contains 12 blocks, with model dimension 768, feed-forward dimension 3072, and 12 attention heads. Before these features are fed to the transformer encoder, we randomly mask them in time dimension with a probability of 0.8. For the transformer encoder, instead of the simple relative position embedding used in HuBERT, we use a bucket relative position embedding [18], generating different learned embeddings based on the offset between the “key” and the “query” in the vanilla self-attention.
@@ III-B Experimental Setup
All pre-training experiments are conducted using fairseq toolkit  [25]. For GAN-based tokenizer, we mainly follow the training process and hyper-parameters in [16]. Specifically, we first use the unsupervised voice activity detection [26] to remove the silences from speech and then feed the speech into the pre-trained speech model, i.e., wav2vec 2.0-large , to extract high-level speech representations. For unpaired text, we convert the word sequences into phoneme sequences using an off-the-shelf phonemizer [27], which contains a total of 41 monophones, including a silence symbol. The generator consists of a single-layer CNN with batch normalization and the discriminator employs a three-layer CNN that takes either the output of the generator or a one-hot representation of the real phoneme sequence as input.
All pre-training models are trained on 8 V100 32GB GPUs with a gradient accumulation number of 4 for a total of 400k steps. We randomly sample starting positions with a probability of 0.08 and apply a mask to the subsequent 10 time steps. We employ the Adam optimizer [28] with a weight decay of 0.01 and $\beta=(0.9,0.98)$. The learning rate follows a linear increase for the first 32k steps and subsequently decays linearly to 0. The peaking learning rate is 5e-4. Following [2, 1], we use the Connectionist Temporal Classification (CTC) [29] loss for downstream ASR fine-tuning. The target vocabulary includes 26 English characters, a space token, and an apostrophe. We train the model with Adam optimizer and a tri-stage rate schedule where the learning rate is warmed up for the first 10% of the updates, held constant for the next 40%, and then linearly decayed for the remainder. During the inference stage, we use the flashlight [30] beam search decoder with beam size 1500 for a 4-gram Language Model (LM)  [^2] https://www.openslr.org/resources/11/4gram.arpa.gz.
@@ IV-A Main Results
To demonstrate the effectiveness of our proposed T-HuBERT model, we first evaluate it on the standard LibriSpeech test sets. We mainly compare T-HuBERT with three kinds of models, including 1) speech-only pre-training models (i.e., wav2vec 2.0, HuBERT and WavLM), 2) speech-text pre-training models that use paired text data (i.e., PBERT and SpeechLM), 3) speech-text pre-training model that uses unpaired text data (i.e., token2vec). Table I shows the WER results when the model is fine-tuned on 1 hour, 10 hours, 100 hours, 360 hours, and 500 hours of labeled data, respectively. For token2vec, we only report the results without LM, since it did not release the decoding results with LM. Firstly, our T-HuBERT achieves state-of-the-art performance with significant improvement overcall methods on low-resource scenarios (i.e., 1-hour and 10-hour subsets). Secondly, on the train-clean-100 split, T-HuBERT obtains 4.8/10.0% WER, which outperforms all pre-trained models that only use speech (i.e., 5.7/12.0% of WavLM) and unpaired speech and text data (i.e., 5.1/11.8% of token2vec). This indicates that our T-HuBERT can better leverage additional text data in an unsupervised manner than token2vec. However, our T-HuBERT (3.1/7.0%) falls behind SpeechLM (2.7/6.2%) on the train-clean-100 subset, which is due to that SpeechLM uses 100 hours of paired data for pre-training, while T-HuBERT does not involve any paired data during the whole training process. To validate the effectiveness of T-HuBERT on the larger datasets, we further conduct fine-tuning experiments using train-clean-360 and train-other-500. The results presented at the bottom of Table I show that our proposed T-HuBERT also outperforms the baseline HuBERT by a large margin.

### v5

@@ II-B GAN-based Tokenizer
To leverage additional text data and improve the phonetic information of the pre-training, we utilize the GANs [17] as a bridge between unpaired speech and text, to discretize the speech waveform into phoneme-like pseudo-label sequences in an unsupervised manner. Specifically, as shown in the right part of Figure 1, the generator consumes speech, while the input of the discriminator is the real phoneme sequences of unpaired text. During adversarial training, the discriminator is trained to distinguish the output of the generator and real phoneme sequences. The goal of the generator is to output phoneme sequences that cannot be distinguished by the discriminator.
@@ II-C Multi-Unit and Multi-Layer Supervision Objective
Our model architecture design mainly follows the HuBERT-base model, consisting of a convolutional feature encoder and a transformer encoder. The convolutional encoder is composed of seven temporal convolution blocks, which have 512 channels with strides $(5,2,2,2,2,2,2)$ and kernel widths $(10,3,3,3,3,2,2)$, resulting in a 20ms frame rate for audio sampled at 16kHz. The transformer encoder contains 12 blocks, with model dimension 768, feed-forward dimension 3072, and 12 attention heads. Before these features are fed to the transformer encoder, we randomly mask them in time dimension with a probability of 0.8. For the transformer encoder, instead of the simple relative position embedding used in HuBERT, we use a bucket relative position embedding [18], which generates different learned embeddings based on the offset between the “key” and the “query” in the vanilla self-attention.
@@ III-B Experimental Setup
All pre-training experiments are conducted using fairseq toolkit  [25]. For GAN-based tokenizer, we mainly follow the training process and hyper-parameters in [16]. Specifically, we first use the unsupervised voice activity detection [26] to remove the silences from speech and then feed the speech into the pre-trained speech model, i.e., wav2vec 2.0-large [^2] https://dl.fbaipublicfiles.com/fairseq/wav2vec/wav2vec_vox_new.pt, to extract high-level speech representations. For unpaired text, we convert the word sequences into phoneme sequences using an off-the-shelf phonemizer [27], which contains a total of 41 monophones, including a silence symbol. The generator consists of a single-layer CNN with batch normalization and the discriminator employs a three-layer CNN that takes either the output of the generator or a one-hot representation of the real phoneme sequence as input.
All pre-training models are trained on 8 V100 32GB GPUs with a gradient accumulation number of 4 for a total of 400k steps.We randomly sample starting positions with a probability of 0.08 and apply a mask to the subsequent 10 time steps. We employ the Adam optimizer [28] with a weight decay of 0.01 and $\beta=(0.9,0.98)$. The learning rate follows a linear increase for the first 32k steps and subsequently decays linearly to 0. The peaking learning rate is 5e-4. Following [2, 1], we use the Connectionist Temporal Classification (CTC) [29] loss for downstream ASR fine-tuning. The target vocabulary includes 26 English characters, a space token, and an apostrophe. We train the model with Adam optimizer and a tri-stage rate schedule where the learning rate is warmed up for the first 10% of the updates, held constant for the next 40%, and then linearly decayed for the remainder. During the inference stage, we use the flashlight [30]  [^3] https://github.com/flashlight/text beam search decoder with beam size 1500 for a 4-gram language model (LM)  [^4] https://www.openslr.org/resources/11/4gram.arpa.gz.
@@ IV-A Main Results
To demonstrate the effectiveness of our proposed T-HuBERT model, we first evaluate it on the standard LibriSpeech test sets. We mainly compare T-HuBERT with three kinds of models, including 1) speech-only pre-training models (i.e., wav2vec 2.0, HuBERT and WavLM), 2) speech-text pre-training models that use paired text data (i.e., PBERT and SpeechLM), 3) speech-text pre-training model that uses unpaired text data (i.e., token2vec). Table I shows the WER results when the model is fine-tuned on 1 hour, 10 hours, 100 hours, 360 hours, and 500 hours of labeled data, respectively. For token2vec, we only report the results without LM, since it did not release the decoding results with LM.Firstly, our T-HuBERT achieves state-of-the-art performance with significant improvement overcall methods on low-resource scenarios (i.e., 1-hour and 10-hour subsets). Secondly, on the train-clean-100 split, T-HuBERT obtains 4.8/10.0% WER, which outperforms all pre-trained models that only use speech (i.e., 5.7/12.0% of WavLM) and unpaired speech and text data (i.e., 5.1/11.8% of token2vec). This indicates that our T-HuBERT can better leverage additional text data in an unsupervised manner than token2vec. However, our T-HuBERT (3.1/7.0%) falls behind SpeechLM (2.7/6.2%) on the train-clean-100 subset, which is due to that SpeechLM uses 100 hours of paired data for pre-training, while T-HuBERT does not involve any paired data during the whole training process.
To validate the effectiveness of T-HuBERT on the larger datasets, we further conduct fine-tuning experiments using train-clean-360 and train-other-500. The results presented at the bottom of Table I show that our proposed T-HuBERT also outperforms the baseline HuBERT by a large margin.